OpenStudy (anonymous):

three points a , b, c are on a straight line such that ab=bc .a particle moving with uniform acceleration a and c with speeds of 6m/s and 8m/s another particle moving with uniform acceleration passes a and b with speeds of 12m/s and 14m/s it t1 and t2 are taken by the particle to cross ab t/t2 is??

OpenStudy (experimentx):

can you clarify the last statement??

OpenStudy (anonymous):

u have find t1/t2 (ratio)

OpenStudy (experimentx):

"it t1 and t2 are taken by the particle to cross ab t/t2 is??" ???

OpenStudy (anonymous):

sorry it is t1/t2 (ratio)

OpenStudy (experimentx):

same as before, use v=u+at

OpenStudy (experimentx):

Oo ... something is wrong!!!

OpenStudy (anonymous):

can u give the full equation

OpenStudy (experimentx):

is AB for both particles??

OpenStudy (anonymous):

yes

OpenStudy (experimentx):

speed of particle 1 at b = (6+8)/2 = 7 now use v = u + at for both ... and take ratios!!

OpenStudy (anonymous):

can u give the solved answer plzzzz

OpenStudy (experimentx):

\( 8 = 7 + at_1 \) \( 14 = 12 + at_2 \)

OpenStudy (anonymous):

thnxxxxx

OpenStudy (experimentx):

didi it work out?? I still have many doubts!!!

OpenStudy (anonymous):

s ur answer is correct

OpenStudy (experimentx):

well ... don't blame if it gets incorrect again ... LOL